Circular No. 6262

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995am IN ANONYMOUS GALAXY M. Lovas, Konkoly Observatory, Budapest, reports his discovery, on

two 103a-D plates taken at Piszkéstető on Oct. 22.832 UT, of a supernova (mag ∼ 15.0) located 112′′ [sic] west and 10′′.0 south of the center of a galaxy located at α = 0^{h}47^{m}50^{s}.14, δ = +29^{o}57′34′′.8 (equinox 2000.0).

P. Garnavich, P. Challis, and R. Kirshner, Center for Astrophysics, write: “A CCD image obtained by A. Goodman on Nov. 17.2 UT with the Mt. Hopkins 1.2-m telescope shows SN 1995am (V ≈ 17.5) as being 9′′ west and 8′′ south of the galaxy center (and suggesting that Lovas’ offset must be in error). A spectrum taken by E. Barton, also on Nov. 17, with the 1.5-m Tillinghast telescope indicates that this is a type-Ia supernova observed 4–6 weeks past maximum. The spectrum of the host galaxy shows narrow emission lines of Hα, [N II], and [S II] at a redshift of 0.017.”

SUPERNOVA 1995an IN UGC 3188 J. Mueller reports her discovery of a supernova (blue mag ∼ 17.5) on

a plate taken with the 1.2-m Oschin Schmidt telescope on Oct. 27 UT by D. Griffith and herself in the course of the second Palomar Sky Survey. SN 1995an is located 11′′ west and 3′′ north of the center of UGC 3188 (α = 4^{h}51^{m}.8, δ = +8^{o}50′, equinox 2000.0), clearly on a spiral arm.

L. Wang and J. C. Wheeler, University of Texas at Austin, comunicate: “A spectrum with low signal-to-noise ratio was obtained on Nov. 17 UT with the 2.1-m telescope at McDonald Observatory; SN 1995an shows prominent Hα emission (FWHM ∼ 8450 km/s) with a weak P-Cyg profile, indicating a type-II supernova. The redshift of the host galaxy is ∼ 3430 km/s as measured from neighboring H II regions.”

COMET 29P/SCHWASSMANN-WACHMANN 1 L. Jorda and J. Lecacheux, Observatoire de Paris, Meudon; and F.

Colas, Bureau des Longitudes, report: “We have imaged this comet on Nov. 14.21 UT with the 1.05-m telescope (+ broadband Schott KG1 filter) at Pic du Midi Observatory. The comet shows a dust shell spanning p.a. 105^{o}–290^{o}, and a prominent jet of length 24′′ (110 000 km at the comet) in p.a. 10^{o}. We estimate m_{1} = 13.7 ± 0.5, from integration within a 1′- diameter field-of-view (and comparison with 3 nearby GSC stars). The

image is available at http://www.bdl.fr/imagespic.html.”

1995 November 17 Daniel W. E. Green